A Gallup poll conducted in 2013 examined whether there is an association between church attendance and smoking status among American adults. The following table reports the percentage of people who classify themselves as smokers, conditional on self-described level of church attendance At least once per week Almost every week About once a month Seldom Never 12 13 23 26 26 Round your answer to 2 decimal places, e.g. 0.83. People who never attend church are weekly to be smokers the absolute tolerance is +/-0.01 times as likely as those who attend church Round your answers to 2 decimal places, e.g. 0.83. People who attend church almost every week are church weekly to be smokers. times as likely as those who attend People who attend church about once a month are church weekly to be smokers. times as likely as those who attend People who attend church seldom are weekly to be smokers. times as likely as those who attend church
Answer #1

(1) 26/12 = 2.17 times

(2) 13/12 = 1.08 times

(3) 23/12 = 1.92 times

(4) 26/12 = 2.17 times